OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? help please

OpenStudy (anonymous):

m = (y1-y)/(x1-x)

OpenStudy (anonymous):

explain ?

OpenStudy (jojo4eva):

-7-3/4-(-1)=-10/5=-2 so the slope is -2

OpenStudy (anonymous):

ohh

OpenStudy (jojo4eva):

you got it??

OpenStudy (anonymous):

yes thank you

OpenStudy (jojo4eva):

alright no problem